[Hallc_running] [New Logentry] RC daily update

simona at jlab.org simona at jlab.org
Sun Feb 2 16:30:01 EST 2020

Logentry Text:
We have been getting decent enough beam since yesterday up until 8 am this morning when Arc 8 box supply faulted: https://logbooks.jlab.org/entry/3776097 , https://cebaf.jlab.org/dtm/open-events?event_id=6035. 

As of now the prognosis for beam restoration to C is sometime during owl. The swing shift leader is on standby and the swing shift TO will be in the counting house as usual. Both Owl shift takers should show up for their shift as usual.

I checked with Steve Lassiter on whether he would want to work on the SHMS Q3 valve during this longer down but he decided not to (tech and cryo support is needed for the work and that's hard to round up on short notice in the weekend). He would like to have the work scheduled as soon as possible though. 

I checked with the target group about the installation of the EPR fiber bundle and I was told that it's better to schedule that work at the same time with the target cell replacement work.

Skyla10 will be rebooted during a regular work-day down when magnet folk abound at the lab. 

Hall C stays in beam permit during this down. The target ladder has been moved to "no target".

The Hall C spin dance is still scheduled for tomorrow morning.

For a status of data taking until this morning when beam left us see: https://logbooks.jlab.org/entry/3776012

No RC meeting today.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3776121
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200202/116db19b/attachment.html>

More information about the Hallc_running mailing list